a)     Suppose the above named Business Company requires an investment of $200, 000 and it is projected to generate an annual cash inflow of $50, 000. Calculate the payback period.

b)     The named international business projects costs initially $25, 000 and generates yearend cash inflows of $9, 000; $7, 000; $6, 000 and $5, 000 from one year to five years. If the required rate of return is 10%, calculate the Net Present Value (NPV).

c)     If the company further projects costs $16, 200 and is expected to generate cash of $8, 000; $7, 000 and $6, 000 over a three year period. Calculate the company’s Internal Rate of Return (IRR).



1
Expert's answer
2021-03-23T11:49:54-0400

a) If Business Company requires an investment of $200,000 and it is projected to generate an annual cash inflow of $50,000, then the payback period is 200,000/50,000 = 4 years.

b) If the international business projects costs initially $25,000, generates year end cash inflows of $9,000; $7,000; $6,000 and $5,000 from one year to five years, and the required rate of return is 10%, then the Net Present Value is:

"NPV = -25,000 + 9,000\/1.1 + 7,000\/1.1^2 + 6,000\/1.1^3 + 5,000\/1.1^4 = -3,110.1."

c) If the company further projects costs $16,200 and is expected to generate cash of $8,000; $7,000 and $6,000 over a three year period, then the company’s Internal Rate of Return (IRR) at which NPV = 0 is:

IRR = 15%.